# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ 4.2.2 Special case with one possible risk aversion
If the distribution of the risk aversion simplifies to a single outcome $\gamma$, then
$\frac{\partial g^{\gamma}(t)}{\partial t}=\frac{1-\gamma}{\gamma}\left[-A-\beta-\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}-\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right]g^{\gamma}(t),$
which yields
$g^{\gamma}(t)=\exp\left\{\frac{1-\gamma}{\gamma}\left(A+\beta+\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}+\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right)(T-t)\right\}.$
Moreover, the equation (4.43) can be rewritten as
@@
In the situation where there is only one possible risk aversion $\gamma\in(0,\epsilon_{2}]$ with $0<\epsilon_{2}<1$, assume that the utility function $\varphi^{\gamma}$ is given by the equation (4.32), $\eta=0$ and the condition $(A2)$ holds. Then the equilibrium reinsurance and investment strategy and the corresponding equilibrium value function are given by equations (4.44) and (4.45), respectively.
@@
If $\alpha=h=B=C=\beta=0$, which suggests that the model does not incorporate any delay, then the equilibrium reinsurance and investment strategy becomes
$\hat{q}(t)=\frac{a\eta_{2}x}{b^{2}\gamma},\quad\hat{\pi}(t)=\frac{(\mu-r)}{\sigma^{2}\gamma}$
and the corresponding equilibrium value function is $V(t,x)=x\exp\left\{\left(r+\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}+\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right)(T-t)\right\}$.
@@ 5 Numerical illustrations
In this section, we will present numerical experiments to analyze the impacts of delay parameters and risk aversion parameters on the equilibrium reinsurance and investment strategies under the exponential utility function and power utility function, respectively. Meanwhile, we will analyze the influences of parameters in the insurance and financial markets on the equilibrium reinsurance and investment strategy under the power utility function. We would like to mention that in the sequel, the sensitivity of the equilibrium reinsurance and investment strategy will be investigated by varying one parameter at a time in each figure and without loss of generality, all investigations will be performed at the initial time $t=0$ for convenience.
@@ 5.1 The case of the exponential utility
Figure 5.1: The impacts of $\alpha$, $\beta$ and $h$ on the equilibrium reinsurance and investment strategy.
Figure 5.1 reflects the impacts of $\alpha$, $\beta$ and $h$ on the equilibrium strategy for reinsurance and investment. As seen in subgraphs 5.1 $(a)$ and $(d)$, both $\hat{q}(0)$ and $\hat{\pi}(0)$ first decrease monotonically as $\alpha$ increases and then increase monotonically. This shows that prior to the extreme point, a higher average parameter corresponds to a smaller reinsurance strategy and a smaller investment proportion in the risky asset, which means the insurer adopts a more conservative reinsurance and investment strategy. Subsequent to the extreme point, a higher average parameter is associated with a greater reinsurance strategy and a greater investment proportion in the risky asset, which implies the insurer uses a more aggressive reinsurance and investment strategy. On the whole, the influence exerted by the average parameter $\alpha$ on the equilibrium reinsurance and investment strategy undergoes a transformation contingent upon the magnitude of the parameter. Moreover, subgraphs 5.1 $(b)$ and $(e)$ disclose that both $\hat{q}(0)$ and $\hat{\pi}(0)$ exhibit a decreasing trend concomitant with the increase of $\beta$. With the increment of $\beta$, the cumulative delayed information related to the insurer’s wealth assumes a greater weight whereas the current wealth carries a smaller proportion. This suggests that the more emphasis the insurer places on the cumulative delayed information of the insurer’s wealth under the exponential utility, the smaller the reinsurance and the proportion to the risky asset will be. Furthermore, subgraphs 5.1 $(c)$ and $(f)$ present that the quantities $\hat{q}(0)$ and $\hat{\pi}(0)$ undergo a reduction in tandem with the growth of $h$. As $h$ increases, the averaging horizon extends, thereby rendering the cumulative delayed information of the wealth more stable. Then, the insurer adopting the exponential utility accepts to reduce the proportion allocated to the risky asset and lower the reinsurance ratio. All aforementioned behaviors correspond to the results in equations (4.28) and (4.29). Additionally, it is readily apparent that a higher expected risk aversion degree for the insurer corresponds to a more cautious reinsurance and investment strategy. Specifically, the insurer with greater expected risk aversion is inclined to acquire additional reinsurance and allocate a smaller proportion to the risky asset. It is also intriguing to note that this phenomenon persists in the subsequent Figures 5.2-5.6 under varying parameter configurations.
@@ 5.2 The case of the power utility
In this subsection, we take $\eta_{1}=\eta_{2}=0.3$ and carry out numerical test to exhibit how delay parameters and insurance-financial market parameters influence $(\hat{q}(0),\hat{\pi}(0))$ under the power utility through Figures 5.2-5.6.
Figures 5.2 and 5.3 demonstrate the impacts of delay parameters $\alpha$, $\beta$ and $h$ on $(\hat{q}(0),\hat{\pi}(0))$. Specifically, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease as $\alpha$ increases, and increase as $\beta$ or $h$ increases. First of all, we can see from the expression of $M_{1}^{u}(t)$ that with an increase in $\alpha$, the weight of earlier wealth in the insurer’s average delayed wealth diminishes, which reflects the insurer attach more importance to wealth nearer to the present time as $\alpha$ grows. Under the power utility function, the insurer would like to procure additional reinsurance and decrease the share of the risky asset in her/his portfolios. What is more, in accordance with the definition of $X^{u}(T)+\beta M_{1}^{u}(T)$, as the delay weight $\beta$ rises, the cumulative delayed information of the insurer’s wealth carries more weight, while the current wealth’s weight is reduced proportionally. Then, the insurer employing the power utility function is more inclined to reduce her/his demand for reinsurance and increase the proportion allocated to the risky asset in portfolios. Additionally, since the delay time $h$ grows, the span of the averaging period lengthens, which in turn stabilizes the cumulative delayed information regarding the insurer’s wealth. This may motivate the insurer who uses the power utility function to decrease the demand for reinsurance and raise the share of the risky asset in investments.
A comparison of Figures 5.1-5.3 reveals that under the exponential utility and power utility, respectively, the same delay parameter changes affect the equilibrium reinsurance and investment strategy differently. This implies that the effects of delay on reinsurance and investment decisions are diverse, thereby reflecting the importance of considering delays.
@@
Figures 5.4 and 5.5 display the effects of insurance market parameters $\mu_{1}$, $\mu_{2}$ and $\eta_{2}$ on $(\hat{q}(0),\hat{\pi}(0))$. Subgraph 5.4 $(a)$ and subgraphs 5.5 $(a)$ and $(d)$ illustrate that both $\hat{q}(0)$ and $\hat{\pi}(0)$ increase monotonically with $\mu_{1}$. We know that $\mu_{1}$ represents the first moment of the claim size. If $\mu_{1}$ increases, then it suggests that the insurer is required to pay a higher premium to the reinsurer, thus encouraging she/he to assume more insurance risks independently and correspondingly reduce the purchase of reinsurance. Under the premise that the insurer is willing to assume more risks, she/he also tends to direct a larger share of investments toward the risky asset to pursue higher returns. Moreover, subgraph 5.4 $(b)$ and subgraphs 5.5 $(b)$ and $(e)$ show that both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease monotonically with $\mu_{2}$. Since $\mu_{2}$ represents the second moment of the claim size, an increase in $\mu_{2}$ implies higher uncertainty within the insurance market. Consequently, the insurer faces higher risks in the insurance market, leading her/him to prefer purchasing more reinsurance to mitigate her/his own risks. Under the premise that the insurer seeks to avoid more risks, she/he is also willing to allocate a smaller proportion of investments to the risky asset. In addition, it can be seen from subgraph 5.4 $(c)$ and subgraphs 5.5 $(c)$ and $(f)$ that both $\hat{q}(0)$ and $\hat{\pi}(0)$ increase alongside a rise in $\eta_{2}$. If the reinsurer’s safety loading $\eta_{2}$ increases, then it implies that the insurer needs to bear higher reinsurance costs. In this case, the insurer becomes more disposed to assume risks by her/his own accord and then reduces the purchase of reinsurance. When the insurer opts to shoulder more risks, she/he also exhibits a tendency to allocate a larger share of investment to the risky asset proportionally.
@@
Figure 5.6 characterizes how financial market parameters $r$, $\mu$ and $\sigma$ affect $(\hat{q}(0),\hat{\pi}(0))$. It follows from subgraph 5.6 $(a)$ that an increase in $r$ leads to a reduction in $\hat{q}(0)$, suggesting that the insurer needs to secure more reinsurance when $r$ rises. Furthermore, subgraph 5.6 $(d)$ demonstrates that $r$ has a similar influence pattern on $\hat{\pi}(0)$. When the risk-free interest rate $r$ increases, returns from risk-free asset investments increase, thereby attracting the insurer to increase the investment proportion in such a asset. Correspondingly, the proportion directed toward the risky asset will decrease. In addition, subgraphs 5.6 $(b)$ and $(e)$ reveal that both $\hat{q}(0)$ and $\hat{\pi}(0)$ rise as $\mu$ increases. As $\mu$ signifies the expected return of the risky asset, an upward movement in $\mu$ indicates higher expected payoffs. Therefore, the insurer is prepared to accept greater risks to expand the share of the risky asset in her/his investments, seeking to gain higher returns. In such a case where the insurer opts to take on more risks, it tends to reduce the reinsurance purchases. What is more, subgraphs 5.6 $(c)$ and $(f)$ show that both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease monotonically with $\sigma$. A rise in the volatility $\sigma$ of the risky asset leads to significant price fluctuations and instability. To avoid risks, insurers will cut the proportion of her/his investments in such a asset. Then, given that the insurer wants to avoid risks, she/he will also increase the purchase of reinsurance.
By comparing Figures 5.4, 5.5 and 5.6, we find that parameters $\mu_{1}$, $\mu_{2}$ and $\eta_{2}$ have more significant impacts on $\hat{q}(0)$, while their impacts on $\hat{\pi}(0)$ are relatively smaller. On the other hand, it can also be seen that parameters $r$, $\mu$ and $\sigma$ have more pronounced effects on $\hat{\pi}(0)$, whereas their effects on $\hat{q}(0)$ are comparatively minor.
@@ 6 Conclusions
This paper focused on the reinsurance and investment optimization issue with delay for an insurer adopting random risk aversion. Specifically, based on a diffusion approximation model for the insurer’s surplus evolution, the insurer is endowed with the flexibility to transfer risks via proportional reinsurance, grow the business through new policy acquisitions, and engage in investments within a Black-Scholes financial market. To begin with, we posed the delayed optimal reinsurance and investment problem with random risk aversion by means of expected certainty equivalent and then formulated the verification theorem to describe the equilibrium reinsurance and investment strategy and the associated equilibrium value function under the game theory framework. Next, we obtained (semi-)analytical equilibrium reinsurance and investment strategies and equilibrium value functions under exponential and power utility frameworks, respectively. Finally, we performed numerical simulations to examine the effects of delay parameters and insurance-financial market parameters on the equilibrium reinsurance and investment strategy. The primary findings are outlined below. (I) A higher expected risk aversion level for the insurer leads to a tendency to buy more reinsurance and assign a smaller investment proportion to the risky asset. (II) The same magnitude of delay parameter changes yields different effects on equilibrium reinsurance and investment strategies for exponential utility versus power utility. To be specific, in the context of exponential utility, both $\hat{q}(0)$ and $\hat{\pi}(0)$ first decrease and then increase when $\alpha$ increases, and decrease when $\beta$ or $h$ increases. In the case of power utility, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease when $\alpha$ increases, and increase when $\beta$ or $h$ increases. (III) Under the power utility, parameters in insurance and financial markets demonstrate analogous influence patterns on the variations of equilibrium reinsurance and investment strategies. In particular, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease when $\mu_{2}$, $r$ or $\sigma$ increases, and increase when $\mu_{1}$, $\eta_{2}$ or $\mu$ increases.
@@
where $K>0$ is a constant. Thus, $M_{1}^{u}(\cdot),M_{2}^{u}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ and then $M_{1}^{u}(\cdot),M_{2}^{u}(\cdot)\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. ∎
@@
The result can be obtained similarly to the proof of Lemma 2.1 in [19], and so we omit the details here. ∎
@@
Our proof methodology adapts the techniques developed in Theorem 15.1 of [10], Theorem 3.5 of [17] and Theorem 3.1 of [25]. To maintain the paper’s completeness, we present the argument in three sequential steps.
Step 1. Given the assumed condition $U,Y^{\gamma},H\in C^{1,2,1}(\mathcal{D})$ for all $\gamma$, it follows that $\hat{u}$ is continuous, which consequently ensures the continuity of both the drift and diffusion coefficients in the SDE governing $X^{\hat{u}}$. Furthermore, given that $Y^{\gamma}$ meets the condition $(A1)$ for all $\gamma$, one can use the Feynman-Kac therorem [38] and equations (3.7) and (3.10) to demonstrate that
$Y^{\gamma}(t,x,m_{1})=\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{\hat{u}}(T)+\beta M_{1}^{\hat{u}}(T)\right)\right]=y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\;\text{for all}\;\gamma.$ (C.46)
@@
$\mathcal{A}^{\hat{u}}U(t,x,m_{1})-\mathcal{A}^{\hat{u}}H(t,x,m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,m_{1}))\mathcal{A}^{\hat{u}}Y^{\gamma}(t,x,m_{1})\mathrm{d}\Gamma(\gamma)=0.$
@@
$\mathcal{A}^{\hat{u}}U(t,x,m_{1})-\mathcal{A}^{\hat{u}}H(t,x,m_{1})=0.$ (C.47)
When $U\in C^{1,2,1}(\mathcal{D})$ and $U(t,X^{u}(t),M_{1}^{u}(t))$ meets the condition $(A1)$, one can apply Lemma 3.2 to obtain that
$\mathbb{E}_{t}\left[U(T,X^{\hat{u}}(T),M_{1}^{\hat{u}}(T))\right]=$ $U(t,x,m_{1})+\mathbb{E}_{t}\left[\int^{T}_{t}\bigg(U_{s}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))+U_{x}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))\right$.
@@
$+0.5\left(b^{2}\hat{q}^{2}(s)+\hat{\pi}^{2}(s)(X^{\hat{u}}(s))^{2}\sigma^{2}\right)U_{xx}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))$
$+\left(X^{\hat{u}}(s)-\alpha M_{1}^{\hat{u}}(s)-e^{-\alpha h}M_{2}^{\hat{u}}(s)\right)U_{m_{1}}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))\bigg)\mathrm{d}s\bigg].$ (C.48)
Similarly, for $H\in C^{1,2,1}(\mathcal{D})$ with $H(t,X^{u}(t),M_{1}^{u}(t))$ satisfying the condition $(A1)$, we derive
$\mathbb{E}_{t}\left[H(T,X^{\hat{u}}(T),M_{1}^{\hat{u}}(T))\right]=$ $H(t,x,m_{1})+\mathbb{E}_{t}\left[\int^{T}_{t}\bigg(H_{s}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))+H_{x}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))\right$.

### v2

@@ 5.2.2 Special case with one possible risk aversion
If the distribution of the risk aversion simplifies to a single outcome $\gamma$, then one can derive that $\frac{\partial g^{\gamma}(t)}{\partial t}=\frac{1-\gamma}{\gamma}\left[-A-\beta-\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}-\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right]g^{\gamma}(t),$ which yields $g^{\gamma}(t)=\exp\left\{\frac{1-\gamma}{\gamma}\left(A+\beta+\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}+\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right)(T-t)\right\}.$ Moreover, the equation (5.53) can be rewritten as
@@
In the situation where there is only one possible risk aversion $\gamma\in(0,\epsilon_{2}]$, assume that the utility function $\varphi^{\gamma}$ is given by the equation (5.42), $\eta=0$ and the condition $(A2)$ holds. Then the equilibrium reinsurance and investment strategy and the corresponding equilibrium value function are given by equations (5.54) and (5.55), respectively.
@@
If $\alpha=h=B=C=\beta=0$, which suggests that the model does not incorporate any delay, then the equilibrium reinsurance and investment strategy becomes $\left(\hat{q}(t),\hat{\pi}(t)\right)=\left(\frac{a\eta_{2}x}{b^{2}\gamma},\frac{(\mu-r)}{\sigma^{2}\gamma}\right)$ and the corresponding equilibrium value function is $V(t,x)=x\exp\left\{\left(r+\frac{0.5(\mu-r)^{2}}{\sigma^{2}\gamma}+\frac{0.5a^{2}\eta^{2}_{2}}{b^{2}\gamma}\right)(T-t)\right\}$.
@@ 6 Numerical illustrations
In this section, we will present numerical experiments to analyze the impacts of delay parameters and risk aversion parameters on the equilibrium reinsurance and investment strategies under the exponential utility function and power utility function, respectively. Meanwhile, under the Black-Scholes model, we will analyze the influences of parameters in the insurance and financial markets on the equilibrium reinsurance and investment strategy under the power utility function. In the sequel, the sensitivity of the equilibrium reinsurance and investment strategy will be investigated by varying one parameter at a time in each figure and without loss of generality, all investigations will be performed at the initial time $t=0$ for convenience.
@@ 6.1 The case of the exponential utility
Figure 6.1: The impacts of $\alpha$, $\beta$ and $h$ on the equilibrium reinsurance strategy.
(a) The impact of $\alpha$ on $\hat{\pi}(0)$
(b) The impact of $\beta$ on $\hat{\pi}(0)$
(c) The impact of $h$ on $\hat{\pi}(0)$
@@
Figure 6.2: The impacts of $\alpha$, $\beta$ and $h$ on the equilibrium investment strategy.
In this subsection, we adopt $\bar{s}_{1}=1.2$ and analyze the impacts of $\alpha$, $\beta$ and $h$ on the equilibrium strategy for reinsurance and investment under the exponential utility by Figures 6.1 and 6.2. As seen in subgraphs 6.1 $(a)$ and subgraphs 6.2 $(a)$ and $(d)$, both $\hat{q}(0)$ and $\hat{\pi}(0)$ first decrease monotonically as $\alpha$ increases and then increase monotonically. This shows that prior to the extreme point, a higher average parameter corresponds to a smaller reinsurance strategy and a smaller investment proportion in the risky asset, which means the insurer adopts a more conservative reinsurance and investment strategy. Subsequent to the extreme point, a higher average parameter is associated with a greater reinsurance strategy and a greater investment proportion in the risky asset, which implies the insurer uses a more aggressive reinsurance and investment strategy. On the whole, the influence exerted by the average parameter $\alpha$ on the equilibrium reinsurance and investment strategy undergoes a transformation contingent upon the magnitude of the parameter. Moreover, subgraphs 6.1 $(b)$ and subgraphs 6.2 $(b)$ and $(e)$ disclose that both $\hat{q}(0)$ and $\hat{\pi}(0)$ exhibit a decreasing trend concomitant with the increase of $\beta$. With the increment of $\beta$, the cumulative delayed information related to the insurer’s wealth assumes a greater weight whereas the current wealth carries a smaller proportion. This suggests that the more emphasis the insurer places on the cumulative delayed information of the insurer’s wealth under the exponential utility, the smaller the reinsurance and the proportion to the risky asset will be. Furthermore, subgraphs 6.1 $(c)$ and subgraphs 6.2 $(c)$ and $(f)$ present that the quantities $\hat{q}(0)$ and $\hat{\pi}(0)$ undergo a reduction in tandem with the growth of $h$. As $h$ increases, the averaging horizon extends, which suggests the insurer adopting the exponential utility accepts to reduce the proportion allocated to the risky asset and lower the reinsurance ratio. All aforementioned behaviors correspond to the results in equations (5.38) and (5.39). What is more, Figure 6.2 shows that $\hat{\pi}(0)$ decreases with the increase of $\delta$. Additionally, it is readily apparent that a higher expected risk aversion degree for the insurer corresponds to a more cautious reinsurance and investment strategy. Specifically, the insurer with greater expected risk aversion is inclined to acquire additional reinsurance and allocate a smaller proportion to the risky asset. It is also intriguing to note that this phenomenon persists in the subsequent Figures 6.3-6.7 under varying parameter configurations.
@@ 6.2 The case of the power utility
Next, we take $\delta=0$ and $\eta_{1}=\eta_{2}=0.3$ and carry out numerical test to exhibit how delay parameters and insurance-financial market parameters influence $(\hat{q}(0),\hat{\pi}(0))$ under the power utility through Figures 6.3-6.7.
@@
Figures 6.3 and 6.4 demonstrate the impacts of delay parameters $\alpha$, $\beta$ and $h$ on $(\hat{q}(0),\hat{\pi}(0))$. Specifically, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease as $\alpha$ increases, and increase as $\beta$ or $h$ increases. First of all, we can see from the expression of $M_{1}^{u}(t)$ that with an increase in $\alpha$, the weight of earlier wealth in the insurer’s average delayed wealth diminishes, which reflects the insurer attach more importance to wealth nearer to the present time as $\alpha$ grows. Under the power utility function, the insurer would like to procure additional reinsurance and decrease the share of the risky asset in her/his portfolios. What is more, in accordance with the definition of $X^{u}(T)+\beta M_{1}^{u}(T)$, as the delay weight $\beta$ rises, the cumulative delayed information of the insurer’s wealth carries more weight, while the current wealth’s weight is reduced proportionally. Then, the insurer employing the power utility function is more inclined to reduce her/his demand for reinsurance and increase the proportion allocated to the risky asset in portfolios. Additionally, since the delay time $h$ grows, the span of the averaging period lengthens, which in turn stabilizes the cumulative delayed information regarding the insurer’s wealth. This may motivate the insurer who uses the power utility function to decrease the demand for reinsurance and raise the share of the risky asset in investments.
@@
A comparison of Figures 6.1-6.4 reveals that under the exponential utility and power utility, respectively, the same delay parameter changes affect the equilibrium reinsurance and investment strategy differently. This implies that the effects of delay on reinsurance and investment decisions are diverse, thereby reflecting the importance of considering delays.
Figures 6.5 and 6.6 display the effects of insurance market parameters $\mu_{1}$, $\mu_{2}$ and $\eta_{2}$ on $(\hat{q}(0),\hat{\pi}(0))$. Subgraph 6.5 $(a)$ and subgraphs 6.6 $(a)$ and $(d)$ illustrate that both $\hat{q}(0)$ and $\hat{\pi}(0)$ increase monotonically with $\mu_{1}$. We know that $\mu_{1}$ represents the first moment of the claim size. If $\mu_{1}$ increases, then it suggests that the insurer is required to pay a higher premium to the reinsurer, thus encouraging she/he to assume more insurance risks independently and correspondingly reduce the purchase of reinsurance. Under the premise that the insurer is willing to assume more risks, she/he also tends to direct a larger share of investments toward the risky asset to pursue higher returns. Moreover, subgraph 6.5 $(b)$ and subgraphs 6.6 $(b)$ and $(e)$ show that both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease monotonically with $\mu_{2}$. Since $\mu_{2}$ represents the second moment of the claim size, an increase in $\mu_{2}$ implies higher uncertainty within the insurance market. Consequently, the insurer faces higher risks in the insurance market, leading her/him to prefer purchasing more reinsurance to mitigate her/his own risks. Under the premise that the insurer seeks to avoid more risks, she/he is also willing to allocate a smaller proportion of investments to the risky asset. In addition, it can be seen from subgraph 6.5 $(c)$ and subgraphs 6.6 $(c)$ and $(f)$ that both $\hat{q}(0)$ and $\hat{\pi}(0)$ increase alongside a rise in $\eta_{2}$. If the reinsurer’s safety loading $\eta_{2}$ increases, then it implies that the insurer needs to bear higher reinsurance costs. In this case, the insurer becomes more disposed to assume risks by her/his own accord and then reduces the purchase of reinsurance. When the insurer opts to shoulder more risks, she/he also exhibits a tendency to allocate a larger share of investment to the risky asset proportionally.
@@
Figure 6.7 characterizes how financial market parameters $r$, $\mu$ and $\sigma$ affect $(\hat{q}(0),\hat{\pi}(0))$. It follows from subgraph 6.7 $(a)$ that an increase in $r$ leads to a reduction in $\hat{q}(0)$, suggesting that the insurer needs to secure more reinsurance when $r$ rises. Furthermore, subgraph 6.7 $(d)$ demonstrates that $r$ has a similar influence pattern on $\hat{\pi}(0)$. When the risk-free interest rate $r$ increases, returns from risk-free asset investments increase, thereby attracting the insurer to increase the investment proportion in such a asset. Correspondingly, the proportion directed toward the risky asset will decrease. In addition, subgraphs 6.7 $(b)$ and $(e)$ reveal that both $\hat{q}(0)$ and $\hat{\pi}(0)$ rise as $\mu$ increases. As $\mu$ signifies the expected return of the risky asset, an upward movement in $\mu$ indicates higher expected payoffs. Therefore, the insurer is prepared to accept greater risks to expand the share of the risky asset in her/his investments, seeking to gain higher returns. In such a case where the insurer opts to take on more risks, it tends to reduce the reinsurance purchases. What is more, subgraphs 6.7 $(c)$ and $(f)$ show that both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease monotonically with $\sigma$. A rise in the volatility $\sigma$ of the risky asset leads to significant price fluctuations and instability. To avoid risks, insurers will cut the proportion of her/his investments in such a asset. Then, given that the insurer wants to avoid risks, she/he will also increase the purchase of reinsurance.
By comparing Figures 6.5, 6.6 and 6.7, we find that parameters $\mu_{1}$, $\mu_{2}$ and $\eta_{2}$ have more significant impacts on $\hat{q}(0)$, while their impacts on $\hat{\pi}(0)$ are relatively smaller. On the other hand, it can also be seen that parameters $r$, $\mu$ and $\sigma$ have more pronounced effects on $\hat{\pi}(0)$, whereas their effects on $\hat{q}(0)$ are comparatively minor.
@@ 7 Conclusions
This paper focused on the reinsurance and investment optimization issue with delay for an insurer adopting random risk aversion. Specifically, based on a diffusion approximation model for the insurer’s surplus evolution, the insurer can transfer risks via proportional reinsurance, grow the business through new policy acquisitions, and engage in investments within the financial market captured by the CEV model. To begin with, we posed the delayed optimal reinsurance and investment problem with random risk aversion by expected certainty equivalent and then formulated the verification theorem to describe the equilibrium reinsurance and investment strategy and the associated equilibrium value function under the game theory framework. Next, we obtained (semi-)analytical equilibrium reinsurance and investment strategies and equilibrium value functions under the CEV model for the exponential utility and under the Black-Scholes model for both exponential and power utilities. Finally, we performed numerical simulations to examine the effects of delay parameters and insurance-financial market parameters on the equilibrium reinsurance and investment strategy. The primary findings are outlined below. (I) A higher expected risk aversion level for the insurer leads to a tendency to buy more reinsurance and assign a smaller investment proportion to the risky asset. (II) The same magnitude of delay parameter changes yields different effects on equilibrium reinsurance and investment strategies for exponential utility versus power utility. To be specific, in the context of exponential utility, both $\hat{q}(0)$ and $\hat{\pi}(0)$ first decrease and then increase when $\alpha$ increases, and decrease when $\beta$ or $h$ increases. In the case of power utility, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease when $\alpha$ increases, and increase when $\beta$ or $h$ increases. (III) Under the exponential utility, the proportion of investment in the risky asset will decrease with the increase of the elasticity parameter $\delta$ if the initial value of the stock price is greater than $1$. (IV) Under the power utility, parameters in insurance and financial markets demonstrate analogous influence patterns on the variations of equilibrium reinsurance and investment strategies. In particular, both $\hat{q}(0)$ and $\hat{\pi}(0)$ decrease when $\mu_{2}$, $r$ or $\sigma$ increases, and increase when $\mu_{1}$, $\eta_{2}$ or $\mu$ increases.
@@
where $K$ is a positive constant that will be used thereafter when there is no ambiguity, which can take different values in different lines. Thus, $M_{1}^{u}(\cdot),M_{2}^{u}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ and then $M_{1}^{u}(\cdot),M_{2}^{u}(\cdot)\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. ∎
@@
The result can be obtained similarly to the proof of Lemma 2.1 in [19], and so we omit details here. ∎
@@
Our proof adapts the techniques developed in Theorem 15.1 of [10], Theorem 3.5 of [17] and Theorem 3.1 of [25]. To maintain the paper’s completeness, we present the argument in three sequential steps.
Step 1. Given the assumed condition $U,Y^{\gamma},H\in C^{1,2,2,1}(\mathcal{D})$ for all $\gamma$, it follows that $\hat{u}$ is continuous, which consequently ensures the continuity of both the drift and diffusion coefficients in the SDE governing $X^{\hat{u}}$. Furthermore, given that $Y^{\gamma}$ meets the condition $(A1)$ for all $\gamma$, one can use the Feynman-Kac therorem [37] and equations (3.7) and (3.10) to demonstrate that
$Y^{\gamma}(t,x,s_{1},m_{1})=\mathbb{E}_{t}\left[\varphi^{\gamma}\left(X^{\hat{u}}(T)+\beta M_{1}^{\hat{u}}(T)\right)\right]=y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\;\text{for all}\;\gamma.$ (C.56)
@@
$\mathcal{A}^{\hat{u}}U(t,x,s_{1},m_{1})-\mathcal{A}^{\hat{u}}H(t,x,s_{1},m_{1})+\int\iota^{\gamma}(Y^{\gamma}(t,x,s_{1},m_{1}))\mathcal{A}^{\hat{u}}Y^{\gamma}(t,x,s_{1},m_{1})\mathrm{d}\Gamma(\gamma)=0.$
@@
$\mathcal{A}^{\hat{u}}U(t,x,s_{1},m_{1})-\mathcal{A}^{\hat{u}}H(t,x,s_{1},m_{1})=0.$ (C.57)
When $U\in C^{1,2,2,1}(\mathcal{D})$ and $U(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))$ meets the condition $(A1)$, one applies Lemma 3.2 to get
$\mathbb{E}_{t}\left[U(T,X^{\hat{u}}(T),S_{1}(T),M_{1}^{\hat{u}}(T))\right]=$ $U(t,x,s_{1},m_{1})+\mathbb{E}_{t}\left[\int^{T}_{t}\bigg(U_{s}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))+U_{x}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))\right$.
@@
$+0.5\left(b^{2}\hat{q}^{2}(s)+\hat{\pi}^{2}(s)(X^{\hat{u}}(s))^{2}\sigma^{2}S^{2\delta}_{1}(s)\right)U_{xx}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\mu S_{1}(s)U_{s_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))+0.5\sigma^{2}S^{2\delta+2}_{1}(s)U_{s_{1}s_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\left(X^{\hat{u}}(s)-\alpha M_{1}^{\hat{u}}(s)-e^{-\alpha h}M_{2}^{\hat{u}}(s)\right)U_{m_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\hat{\pi}(s)X^{\hat{u}}(s)\sigma^{2}S^{2\delta+1}_{1}(s)U_{xs_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))\bigg)\mathrm{d}s\bigg].$ (C.58)
Similarly, for $H\in C^{1,2,2,1}(\mathcal{D})$ with $H(t,X^{u}(t),S_{1}(t),M_{1}^{u}(t))$ satisfying the condition $(A1)$, we derive
$\mathbb{E}_{t}\left[H(T,X^{\hat{u}}(T),S_{1}(T),M_{1}^{\hat{u}}(T))\right]=$ $H(t,x,s_{1},m_{1})+\mathbb{E}_{t}\left[\int^{T}_{t}\bigg(H_{s}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))+H_{x}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))\right$.